(2.1) Find the components of a unit vector satisfying ~v· < 3, −1 >= 0.

(2.2) Show that there are infinitely many vectors in R (4) 3 with Euclidean norm 1 whose Euclidean inner product with < −1, 3, −5 > is zero.

(2.3) Determine all values of k so that ~u =< −3, 2k, −k > is orthogonal to ~v =< 2, (3) 5 2 , −k >


1
Expert's answer
2021-07-11T16:34:09-0400

2.1) Let the vector v = <x, y>


Now v . <3, -1> = 0


<x, y> . <3, -1> = 0

3x - y = 0


One values of x, y satisfying the above equation is (2, 6)


Hence, v = <2, 6>



Now let the unit vector in the direction of v be w.


Therefore,


w = "\dfrac{v}{|v|}" = "\dfrac{2i + 6j}{\sqrt{(2)^2 + (6)^2}}"


w = "\dfrac{2i + 6j}{\sqrt{40}}"





2.2) The given vector is a = <-1, 3, -5>


Let there be a vector v = <x, y, z>


Let the normalized vector of v be V



V = "\dfrac{v}{|v|}"


V = "\dfrac{x}{\sqrt{x^2+y^2+z^2}}, \dfrac{y}{\sqrt{x^2+y^2+z^2}}, \dfrac{z}{\sqrt{x^2+y^2+z^2}}" ....................Equation(1)


Now as per the question, a . V = 0


Hence on taking the dot product we have


-x + 3y - 5z = 0

x = 3y - 5z


On substituting the value of x in equation (1)


V = "\dfrac{3y - 5z}{\sqrt{(3y - 5z)^2+y^2+z^2}}, \dfrac{y}{\sqrt{(3y - 5z)^2+y^2+z^2}}, \dfrac{z}{\sqrt{(3y-5z)^2+y^2+z^2}}"




V is a unitary vector orthogonal to a. Since it still depends on 2 variables we can conclude that there are infinitely many vectors whose inner dot product with a is zero.








2.3) u =< −3, 2k, −k > 


v =< 2, 52 , −k >



Since, u and v are orthogonal so



u . v = 0

<-3, 2k, -k> . <2, 52, -k> = 0


-6 + 104k + k2 = 0


On solving we get


k = 0.0576603

k = −104.058